Energy & Work A elevator has 120505 J of work done on it to ralse it 5.7 m, how much force is needed to lift the elevator (and passengers) if it rising at a constant speed? Answer: 21141.23 N Check Correct Marks for this submission: 1.00/1.00. What is the mass of the elevator (and passengers) if it rising at a constant speed? ofAnswer: 120505 kg Check Incorrect Marks for this submission: 0.00/1.00
Answer #1

Given : W = 120505 J , d = 5.7 m

Constant: Acceleration due to gravity (g)= 9.81 m/s2

Solution:

(1) Force needed to lift the elevator

We know that , work done is given by: W = F*d

Where F is the force needed to lift the elevator

So, 120505 J = F*(5.7 m)

F = 120505 J/5.7 m = 21141.23 N

Answer : F = 21141.23 N

(2) Mass of elevator(and passengers)

The weight of the elevator is given by: w = mg

Where m is the mass of elevator and passengers

i.e. w = mg = 21141.23 N

m = 21141.23 N ÷ 9.81 m/s2 = 2155.07 kg

Answer: m = 2155.07 kg